OpenStudy (anonymous):

Which of the following points is not on the circle represented by the equation (x − 3)2 + (y + 4)2 = 25? Select one of the options below as your answer: A. (6, 0) B. (−1, −1) C. (4, 1) D. None of these is on the circle.

OpenStudy (anonymous):

plug in the (x,y) point in the equation and see if it's true. for the first (x, y) poiny: (6, 0), plugging it in gets: (6 - 3)^2 + (0 + 4)^2 = 25 3^2 + 4^2 = 25 = 9 + 16 = 25 since this is true. this point is on the circle.